OpenStudy (anonymous):

Can someone help me with this inequality? -12x+8 is greater than or equal to 3(-9x+1)

OpenStudy (anonymous):

\[-12x+8 \ge 3(-9x+1)\] 1) multiply out and get variables on one side and numbers on the other.

OpenStudy (anonymous):

\[-12x+8 \ge (-27x+3)\] \[-12x+8 +27x \ge -27x+3 +27x\] you should be able to finish from here:)

OpenStudy (anonymous):

Thanks!